YNWMelly:

What is f(−x) of f(x)=x^3+1 (This is an example I made based off the real question and I really want to know how to solve it)

JammarWalker:

@Hoodmemes

MidNiGhtSkY:

so do y=x^3+1 and find X the opposite of that

jhonyy9:

given f(x) = x^3 +1 to find f(-x) = ? just substitute in place of x the -x in this way f(-x) = (-x)^3 +1 questions ?